OpenStudy (anonymous):

What is the range of the function y = x 2? all real numbers x ≥ 0 y ≥ 0

OpenStudy (anonymous):

y=x2?

OpenStudy (anonymous):

The range is [0,infinity)

OpenStudy (anonymous):

I suppose it is x squared.

OpenStudy (anonymous):

it is squared

OpenStudy (anonymous):

Then what I've written is correct.

OpenStudy (anonymous):

which would it be all real numbers x ≥ 0 y ≥ 0

OpenStudy (anonymous):

All real numbers would be numbers in the range \[\left( -\infty,\infty \right)\]